# Low-Rank Bandits via Tight Two-to-Infinity Singular Subspace Recovery

### v1

@@ I.2 Improved max-norm guarantees
for $\left(i,j\right)\in[m]\times[n]$. We also denote by SBPI the algorithm that outputs the policy $\hat{\pi}_{\text{\tt SBPI}}$ defined by $\hat{\pi}_{\text{\tt SBPI}}\left(i\right)=\arg\max_{1\leq j\leq n}\widehat{M}_{i,j}.$ We finally note that SIPS (resp. SBPI) are denoted by $\operatorname{DSM-PE}$ (resp. $\operatorname{DSM-BPI}$) on the graphs.
We perform numerical experiments on synthetic data with a uniform context distribution. Unless specified otherwise, the behavior policy is uniform, the target policy is chosen as the best policy: $\pi(i)=\arg\max_{j}M_{i,j}$ (ties are broken arbitrarily), and we generate noisy entries $M_{i_{t},j_{t}}+\xi_{t}$ where $\xi_{t}\sim\mathcal{N}\left(0,1\right)$ is standard Gaussian, and where $M=PDQ$ for two invertible matrices $P\in\mathbb{R}^{m\times m}$,$Q\in\mathbb{R}^{n\times n}$ and where $D\in\mathbb{R}^{m\times n}$ is defined by $D_{i,j}=\mathds{1}_{i=j}\mathds{1}_{i\leq r}$ (note that $M$ is consequently of rank $r$). $P$ and $Q$ are initially generated at random with uniform entries in $\left[0,1\right]$ and their diagonal elements are replaced by the sum of the corresponding row to ensure invertibility. All experiments are performed $50$ times and the shaded regions are the corresponding $\left[5\%,95\%\right]$ confidence intervals.
@@
Experiment 1: Impact of data splitting. For a regularization parameter of $\tau=10^{-4}$, we compare the performance of RS-PE for $\alpha\in\left\{1/5,1/2,4/5\right\}$, where $\alpha$ is the proportion of samples used in the first phase of the algorithm. We also plot the error of RS-PE when no data splitting is performed, that is to say when all of the samples are used in both phases of the algorithm. Finally, we plot the dominant term $\frac{\|\psi_{\pi}\|_{2}}{\sqrt{\omega_{\min}}}\sqrt{\frac{\log(16/\delta)}{\left(1-\alpha\right)T}}$ in the instance-dependent upper bound on the PE error of RS-PE$\left(\alpha,10^{-4}\right)$ for $\alpha=4/5$ and $\delta=10^{-2}$ (see Theorem D.7), which approaches its true error bound when $T$ grows large. Note that consequently, the error of RS-PE can be below this dominant term. Additionally, the scaling in $1/\sqrt{1-\alpha}$ of this term suggests that using more samples in the second phase should yield better asymptotic performance. The results are presented in Figure 1.
@@
Surprisingly, all PE estimators appear to have a comparable scaling in $m$. Nonetheless, RS-PE outperforms the two other estimators for every $m$, even though we have chosen a reward matrix for which $r=1$, $\mu=1$, $\kappa=1$ and $\|M\|_{\max}=1$, which is the case where the error guarantee of SIPS matches the RS-PE one up to logarithmic factors.

### v2

@@ I.3 Improved max-norm guarantees
for $\left(i,j\right)\in[m]\times[n]$. We also denote by SBPI the algorithm that outputs the policy $\hat{\pi}_{\text{\tt SBPI}}$ defined by $\hat{\pi}_{\text{\tt SBPI}}\left(i\right)=\arg\max_{1\leq j\leq n}\widehat{M}_{i,j}.$ We finally note that SIPS (resp. SBPI) is denoted by $\operatorname{DSM-PE}$ (resp. $\operatorname{DSM-BPI}$) on the graphs.
We perform numerical experiments on synthetic data with a uniform context distribution. Unless specified otherwise, the behavior policy is uniform, the target policy is chosen as the best policy: $\pi(i)=\arg\max_{j}M_{i,j}$ (ties are broken arbitrarily), and we generate noisy entries $M_{i_{t},j_{t}}+\xi_{t}$ where $\xi_{t}\sim\mathcal{N}\left(0,1\right)$ is standard Gaussian, and where $M=PDQ$ for two invertible matrices $P\in\mathbb{R}^{m\times m}$, $Q\in\mathbb{R}^{n\times n}$, and $D\in\mathbb{R}^{m\times n}$ defined by $D_{i,j}=\mathds{1}_{i=j}\mathds{1}_{i\leq r}$ (note that $M$ is consequently of rank $r$). $P$ and $Q$ are initially generated at random with uniform entries in $\left[0,1\right]$ and their diagonal elements are replaced by the sum of the corresponding row to ensure invertibility. All experiments are performed $50$ times and the shaded regions are the corresponding $\left[5\%,95\%\right]$ confidence intervals [^22] The code used in the experiments can be accessed at https://github.com/wilrev/LowRankBanditsTwoToInfinity..
@@
Experiment 1: Impact of data splitting. For a regularization parameter of $\tau=10^{-4}$, we compare the performance of RS-PE for $\alpha\in\left\{1/5,1/2,4/5\right\}$, where $\alpha$ is the proportion of samples used in the first phase of the algorithm. We also plot the error of RS-PE when no data splitting is performed, that is to say when all of the samples are used in both phases of the algorithm. Finally, we plot the dominant term $\frac{\|\psi_{\pi}\|_{2}}{\sqrt{\omega_{\min}}}\sqrt{\frac{2\log(16/\delta)}{\left(1-\alpha\right)T}}$ in the instance-dependent upper bound on the PE error of RS-PE for $\alpha=4/5$ and $\delta=10^{-2}$ (see Theorem D.7), which approaches its true error bound when $T$ grows large. Note that consequently, the error of RS-PE can be below this dominant term. Additionally, the scaling in $1/\sqrt{1-\alpha}$ of this term suggests that using more samples in the second phase should yield better asymptotic performance. The results are presented in Figure 1.
@@
Surprisingly, all PE estimators appear to have a comparable scaling in $m$. Nonetheless, RS-PE outperforms the two other estimators for every $m$, even though we have chosen a reward matrix for which $\|M\|_{\max}=\mu=\kappa=r=1$, so that the error guarantee of SIPS matches the RS-PE one up to logarithmic factors.